### v1

# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

Abstract
[…]
Uncovering Autoregressive LLM Knowledge of Thematic Fit
in Event Representation
Safeyah Khaled Alshemali Columbia University New York, NY, USA ska2145@columbia.edu                        Daniel Bauer Columbia University New York, NY, USA db2711@columbia.edu                        Yuval Marton University of Washington WA, USA ymarton@uw.edu
1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge (including linguistic knowledge) about the thematic fit of a given event argument to its assigned semantic role for a given predicate (the event ‘trigger’)? A naive approach of prompting LLMs might yield negative results and thus lead to wrong conclusions about the absence of such knowledge. However, we hypothesize that this knowledge does exist, at least to a certain degree, but needs eliciting. We test our hypothesis along three axes: (1) Whether structured reasoning in prompting improves the model’s estimations, following the success of chain-of-thought  prompting (Wei et al., 2022). (2) Whether providing generated sentences as contexts helps, compared to providing the model with only the lemma of the predicate, the lemma of the argument’s syntactic head, and its semantic role. (3) Whether asking for categorical output (such as “low” or “high” thematic fit) yields better results than asking for a numeric output $(0..1)$ as the model’s thematic fit estimation.
Thematic Fit
The thematic fit estimation task measures the compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a specific semantic role assigned to the argument. [^1] Thematic fit is related to, but separate from, semantic role labeling (SRL). Given a sentence, SRL should output the semantic frames in the sentence, including the predicates, arguments, and their semantic roles in the frame. Thematic fit takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (usually denoted as Arg0  in PropBank; see sec.§3.2), then ‘people’, ‘husband’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1  in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. And so on. Psycholinguistic experiments show that humans rely on generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (McRae and Matsuki, 2009; Bicknell et al., 2010). The thematic fit task examines how well computational event representations are aligned with this type of human event knowledge.
The thematic fit task is challenging due to the lack of direct training data. Thematic fit norms (see §3.2), which summarize human judgments, provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, thematic fit models are traditionally trained on proxy data (either PropBank (Kingsbury and Palmer, 2002) or FrameNet (Johnson et al., 2016) annotations). Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (Levy et al., 2017), similarity judgments (Hill et al., 2015), and more. However, on thematic fit, these models, including masked LMs such as BERT (Devlin et al., 2019), have so far only yielded modest improvements over previous approaches (Pedinotti et al., 2021). In this work, we examine if more recent, large autoregressive models, such as GPT (Radford et al., 2018; Achiam et al., 2023), can produce better results using prompting. We attempt to elicit their relevant linguistic knowledge along the following three axes:
Step-by-Step Prompting
Several studies show LLMs perform remarkably well in challenging NLP tasks such as Planning (Wang et al., 2022), automatic scoring (Lee et al., 2024), and question answering (Wang et al., 2023), when using ‘chain-of-thought  prompting’ (Wei et al., 2022). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise. We view thematic fit estimation as a linguistic reasoning task because it can be broken into clear sub-steps: listing the argument’s (salient or relevant) properties, listing the (prototypical or necessary) properties required for the role “assigned” by the given predicate, and estimating the argument’s thematic fit to the assigned role given these two lists. Thus, we compare the LLMs’ performance using a “naive” prompting approach (Simple Prompting hereafter) with the chain-of-thought  approach (Step-by-Step Prompting  hereafter) on the thematic fit estimation task.
Generated Sentences vs. Lemma Tuples
The test sets we use in this study only provide tuples of lemmas, for example, (predicate: eat, argument: pizza, role: Arg0/Agent), and a thematic fit score (an average of several human raters’ scores). Since LLMs were trained on full sentences (as opposed to lemma tuples), we assume this format will be unnatural for them, even if the relevant linguistic and/or world knowledge exists. Therefore, we attempt to reconstruct full sentences with these predicates, arguments, and roles. Furthermore, upon a close reading, we find that the human raters were indeed provided full sentences for the rating (McRae et al., 1998), but this fact seems to have been lost over the years, and only the lemma tuple format has been used (Tilk et al., 2016; Hong et al., 2018; Muthupari et al., 2022). We close here a full circle, bringing the conditions closer to the original.
Predefined Categories Vs. Numeric Score
The thematic fit ratings in our test sets are Likert scale ratings (following McRae et al. (1998)) averaged over 7 human raters. Thus, a straightforward evaluation would prompt the model to output numeric scores as well. For each test set, we then calculate the Spearman Rank Correlation Coefficient (§3.4) LLMs  were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. In fact the LLMs’ numeric scores are actually a textual output, which we then convert to its corresponding numeric value. We, therefore, also prompt the LLMs to score the thematic fit according to predefined categories, which we then convert to corresponding predefined values as well (and then compute a correlation score as before).
Contributions
Our main contributions are:
  • First study to use prompting autoregressive LLMs to directly estimate thematic fit.
  • Comparison of Simple Prompting  vs. Step-by-Step Prompting  for thematic fit estimation.
  • Comparison of generated sentences as additional input context vs. only using lemma forms of the predicate and argument’s syntactic head.
  • Comparison of numeric outputs vs. predefined categories (“Low”, “High”, etc.)
  • New state-of-the-art with our GPT-powered method on all tested tasks (see details of fair comparison to the baselines in Sect. §4).
2 Related Work
Early distributional approaches to the thematic fit task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (Erk et al., 2010; Baroni and Lenci, 2010; Santus et al., 2017; Sayeed et al., 2015; Greenberg et al., 2015). Our first baseline (hereafter BG) is the distributional approach by Greenberg et al., which is based on the syntactically structured distributional semantic model by Baroni and Lenci. More recent approaches use neural networks to learn event representations. Tilk et al. (2016) describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. Hong et al. (2018) improve over Tilk et al. by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1).Marton and Sayeed (2021) replicate Hong et al. (2018)’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). Muthupari et al. (2022) focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the Hong et al. model with random vectors vs. pre-trained GloVe embeddings (Pennington et al., 2014), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
Contemporary work using pretrained LLMs (without any additional supervised training) for thematic fit either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’) (Kauf et al., 2023), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence Metheniti et al. (2020); Pedinotti et al. (2021). Notably, Pedinotti et al. only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (Chersoni et al., 2019). Finally, Testa et al. (2023) examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis. We are the first to use Simple and Step-by-Step Prompting with autoregressive LLMs directly for thematic fit.
3 Experiments
The experiments were designed to investigate three axes with two possible values each. Therefore, we conduct $82^{3}=$ 8 experiments (see Table 1), labeled according to these axes. Each experiment has a prompting template according to the input, output, and reasoning settings, as detailed in Appendix §A.
Reasoning Form Input Form Output Form
Simple Prompting Head Lemma Tuples 1.1 Num. Score
1.2 Cat. Score
Generated Sentences 2.1 Num. Score
2.2 Cat. Score
Step-by-Step Prompting Head Lemma Tuples 3.1 Num. Score
3.2 Cat. Score
Generated Sentences 4.1 Num. Score
4.2 Cat. Score
Table 1: Overview of experiments. ’Num.’ = numeric score. ’Cat.’ = categorical score.
Figure 1: Simple Prompting Approach
Figure 2: Step-by-Step Prompting Approach
3.1 Experiment Design: Three Axes
Axis1: Reasoning Form: Simple Prompting  vs. Step-by-Step Prompting
The first axis is concerned with the effectiveness of Step-by-Step Prompting  (Exp.3.x and 4.x) compared to Simple Prompting  (Exp.1.x and 2.x) in evaluating the thematic fit (leftmost column in Table 1). In the Simple Prompting  approach, the model is asked directly only one prompt, which we call a ‘simple thematic fit prompt’, to score the thematic fit of some predicate’s argument to its assigned semantic role (where the predicate, the argument, and the semantic role are all given).
[…]
Axis 2: Input Form: Head Lemma Tuples vs. Generated Sentences
[…]
Axis 3: Output Form: Categorical vs. Numeric Thematic Fit Scoring
Each experiment may generate one of two forms of output: a numeric thematic fit score (Exp.x.1) or a predefined thematic fit category (Exp.x.2). The numeric thematic fit score (numeric score, hereafter) is a numeric value in the range [0,1]. The predefined thematic fit categories (categorical score, hereafter) are ’Near-Impossible’, ’Low’, ’Medium’, ’High’, or ’Near-Perfect’. These output score categories are then converted to the following numeric values: (0.0, 0.25, 0.5, 0.75, 1.0), respectively, in a post-processing script, in order to compute a correlation score compared to the human ratings.
3.2 Datasets and Evaluation
Training: We used pre-trained LLMs (see sec. §3.3) “off the shelf” without modification or additional training. In an ideal world, we would be able to point the reader to the full pre-training data. Test sets: We used the following datasets to evaluate our methods.
McRae et al. (1998), hereafter ’McRae’, contains $1,444$ predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0  and Arg1 (Kingsbury and Palmer, 2002). We removed eight duplicate records.
Padó et al. (2006), hereafter ’Pado’, contains 414 tuples in the same form as McRae et al. (1998); however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred volunteer English native speakers judged these tuples in four separate randomly ordered lists.
Ferretti et al. (2001) evaluate thematic fit for the roles Instrument  (hereafter, the Fer-Ins dataset) and Location  (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in McRae et al. (1998). However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
Evaluation: We used Spearman’s Rank Correlation (Spearman, 1904) between each of the test sets (each set sorted by the average human ratings for each item) and the corresponding model output (sorted as well). Categorical output was converted to numeric values (detailed in §3.1).
3.3 Models and Hardware
We conducted the experiments using two models, the latest as of the time of running the experiments. (1) OpenAI GPT-4-turbo (gpt-4-0125-preview, hereafter ‘GPT’) (Achiam et al., 2023) as a closed model, and (2) Meta CodeLlama2 (codeLlama2-70B-instruct, hereafter ‘Llama’) (Roziere et al., 2023) as an open model (even if not fully open). See details and required hardware in Appendix §F.
3.4 Experimental Settings
We conducted a preliminary study to determine good temperature and top_p parameter values for the LLMs. We used GPT-3.5 on a subset of the Fer-Ins dataset (20 samples $\sim$ 8%) and ran a parameter sweep over temperature values (0.0, 0.5, 0.9) and top_p values (0.5, 0.7, 0.95). Temperature values were chosen near-uniformly across the scale to examine a wide range of the model creativity setting. Higher values correspond to higher randomness (aka "creativity") in the output. High values of top_p limit the randomness in the output to more probable tokens. Since we were more interested in accurate output, and less interested in creative writing for our purposes here, we tested top-p threshold values on the higher half of the scale. We then sort the human rating list and the generated scores list and compute a Spearman’s Rank Correlation Coefficient ($\rho$) between these two sorted lists. Temperature 0.0 and top_p 0.95 showed the highest correlation coefficient with human judgments on our sample. We set the max_tokens parameter based on the length of the prompt and its expected output. For experiments with Head Lemma Tuples input, we set max_tokens to 100, whereas for Generated Sentences input, we set it to 300. As Step-by-Step Prompting  is expected to be lengthy to clarify the underlying reasons, we set max_tokens to 600 there. We used these settings in all experiments. [^3] The source code is available at: https://github.com/SafeyahShemali/LLM-Thematic-Fit-24
4 Results
This section consists of the following parts. §4.1: the GPT  results. §4.2: the Llama results.
4.1 GPT Results
Experiment 1.1, the simplest in settings and the most similar to the baselines (using head lemmas), yielded scores that are pronouncedly higher than the baselines across the board. [^4] Except for baseline B3c on Pado, whose results are slightly higher than GPT’s. But B3c reports the highest scores in several runs, while the average is reported in B3b, which is slightly lower than the score in Exp.1.1. See Table 2. This is a surprising and impressive achievement for GPT, since to the best of our knowledge, it does not have an explicit self-reflection capability nor was it trained on thematic fit labels.
Axis 1: Reasoning Form
Step-by-Step Prompting  dramatically improved Ferretti’s Location (Fer-Loc) results compared to Simple Prompting, yielding a new state-of-the-art (.31 absolute improvement over the strongest baseline). Results in all Step-by-Step Prompting  experiments (Exp.3.x or 4.x) were higher than all baselines on the Ferretti datasets.
For Fer-Ins, where the highest score with Step-by-Step Prompting was .68, the correlation coefficients improved with Step-by-Step Prompting for .18-.20 (absolute) over the strongest baseline. However, the maximum score on Fer-Ins (.74) was obtained with Simple Prompting  and categorical output form (Exp.1.2).
Surprisingly, Step-by-Step Prompting  achieved much lower scores on Pado and McRae compared with Simple Prompting in Exp.3.x and slightly lowered scores in Exp.4.x. Overall, there was a sizable gain in Fer-Loc, but only there.
Results from Exp.1, 2, and 4 were generally higher than the baseline results. [^5] But for Pado, see footnote 4. In Exp.3.x, the scores on Pado were lower than all of the baselines. Similarly, the scores on McRae were lower than most baselines. However, the corresponding scores in Exp.4.x (Step-by-Step Prompting  with generated sentences) were much higher.
Axis 2: Input Form
Using Generated Sentences instead of Lemma Tuples yielded large gains with Step-by-Step Prompting  on Pado and McRae. The gain in Pado was .26 (absolute) with a numeric score and .25 (absolute) with a categorical score. The gain in McRae was .09 (absolute) with either a numeric or categorical score.
Using Generated Sentences instead of Lemma Tuple with Simple Prompting  lowered thematic fit scores on all test sets except McRae in Exp.2.2 vs. Exp.1.2 (Simple Prompting  and categorical scores), where we got a slight gain, if any, of .03 (absolute).
Axis 3: Output Form
Using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) raised the correlation coefficient across the board in all experiments and test sets, with few exceptions. Most notable of which were the lower scores of Exp.1.2 compared with 1.1 on Pado and McRae. The other case was with Pado in Exp.4.2.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .62 .73 .56 .50
Exp.1.2 .62 .74 .54 .48
Exp.2.1 .43 .63 .52 .48
Exp.2.2 .52 .68 .54 .51
Exp.3.1 .77 .67 .28 .33
Exp.3.2 .77 .68 .28 .36
Exp.4.1 .53 .66 .54 .42
Exp.4.2 .57 .68 .53 .45
Table 2: GPT Spearman Rank Correlation coefficient ($\rho$) results. BG = Greenberg et al. (2015), B0 = Tilk et al. (2016), B1 = Hong et al. (2018), B2 = Marton and Sayeed (2021), B3 = Muthupari et al. (2022) as baselines for our experiments. Subscripts a-c denote specific models there (Glove-shared-not-tuned, etc.). All $\rho$’s had p-values $<10^{-7}$.
4.2 Llama Results
Llama  results on our 8 experiments and 4 datasets are shown in Table 5. In its basic reasoning setting (Exp.1.x), Llama is on par with some of the previous work baselines (e.g., BG) on both Fer-Loc and Fer-Ins, although weaker than other baselines. Interestingly, Llama yielded low-to-random correlation scores on Pado and McRae. Overall, Llama  scores were much lower than those of GPT.
Axis 1: Reasoning Form
Step-by-Step Prompting  (Exp.3.1) yielded much higher scores than the Simple Prompting  (Exp.1.x) on the Ferretti datasets. Its scores on these datasets were higher than all the baselines and would have constituted new state-of-the-art scores, if not for the GPT  scores that were even higher (§4.1). Curiously, this happened only with numeric output but not with categorical output (Exp.3.2), where scores actually drop severely.
Overall, except for Exp.3.1 and Exp.4.1 on the Ferretti datasets, Llama  did not yield better results than any of the baselines.
Axis 2: Input Form
Using Generated Sentences did not improve results across the board, except for: (1) Fer-Loc with categorical output (Exp.2.2 vs. Exp.1.2). (2) Fer-Ins with numerical output (Exp.2.1 vs. Exp.1.1). (3) Pado with numerical output (Exp.2.1 vs. Exp.1.1 and Exp.4.1 vs. Exp.3.1).
Axis 3: Output Form
Using our predefined categorical output yielded lower correlation coefficient compared with numeric output across the board, except for Pado in Exp.3.2 with no or slight gain of .01 (absolute). No x.2 experiments yielded results higher than the previous work baselines, except for Exp.3.2 and Exp.4.2 on Fer-Ins, where the scores (.36-.37) were greater than B3a and B3b.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .31 .42 .08 .22
Exp.1.2 .18 .22 -.06 .00
Exp.2.1 .26 .43 .16 .17
Exp.2.2 .21 .19 -.04 -.01
Exp.3.1 .48 .50 .14 .15
Exp.3.2 .21 .37 .16 .10
Exp.4.1 .39 .46 .20 .13
Exp.4.2 .14 .36 -.01 .09
Table 3: Llama Spearman Rank Correlation coefficient ($\rho$) results. Same models as in Table 2. All $\rho$’s for Fer-Loc and Fer-Ins had p<.003; for Pado and McRae we had p<.84 and p<.94, respectively (very high p-value).
5 Discussion
The core question we aimed to address in this work is: Do LLMs have the kind of linguistic knowledge required for thematic fit estimation? If so, (Q1) does Step-by-Step Prompting  help utilize the model’s internal linguistics knowledge for this task? (Q2) Would providing more context by including sentences as part of the input (as opposed to providing only lemma tuples and semantic roles) improve the model’s thematic fit estimation? (Q3) Do predefined output categories for thematic fit estimation better elicit the model’s internal linguistic knowledge compared to numeric output scores?
Overall, we conclude that GPT  acquired much of the linguistic knowledge required for thematic fit estimation. Results with GPT  on each of the test sets, at least in some of the experimental settings, achieved or even surpassed all previous baselines, which were trained on labeled linguistic data (even if not directly trained for these test sets). This is impressive, assuming GPT  was not trained on these test sets (Sadly, we cannot know for sure because GPT  is a closed model). Next, we discuss the findings by the axes (as defined in §3) and then move to analyzing additional aspects.
5.1 Analysis by our three axes
Axis 1 (Q1): Reasoning Form
On Fer-Loc, Step-by-Step Prompting  fulfilled its intended goal of providing both LLMs  with a “scratchpad” to help them calculate inferences given their internal knowledge – with both models. This held for Fer-Ins too using Llama, but it held only with generated sentences using GPT. Step-by-Step Prompting did not generally help for the Pado and McRae tasks. We hypothesize that interpreting Location  and Instrument  roles is easier for the models, as the role names are self-explanatory, unlike the opaque ProbBank roles. We conjecture that the higher performance on Location relative to Instrument stems from the kind of reasoning required to estimate thematic fit for it, which is simpler for Location than that required for Instrument (and Location  mentions are much more frequent). Whatever reasoning LLMs can do, their current weak level may only suffice for simpler cases. See further analysis in §5.2 (Fer-Loc vs. Fer-Ins).
Why did reasoning have a mixed effect? In further analysis (Appendix §D), we examined the connection between "bad" (invalid, irrelevant, etc.) reasoning and thematic fit scores, using a heuristic to drill down from list-level correlation score to individual items’ fit estimation quality. We concluded that GPT better estimated thematic fit when the reasoning was "good", and tended to perform worse when the reasoning was "bad". We also noticed that in all sampled cases, when "bad" reasoning happened early on in the reasoning chain, it negatively affected the rest of the chain.
Axis 2 (Q2): Input Form
In principle, providing more context enhances models’ ability to comprehend complex tasks. However, this was not generally reflected in our experiments (with the exception of GPT  on Pado and McRae with Step-by-Step Prompting). The reason could be that LLMs  may not link roles like ’Arg1’ with their PropBank sense (more on that in §5.2), or even not properly represent the meaning of roles such as Instrument and Location. This may explain why the LLMs  generated many semantically incoherent sentences (including surface-wise coherent sentences but not with the requested roles), sometimes even up to 50% in small random samples we analyzed.
Axis 3 (Q3): Output Form
We expected predefined output categories to be better than numeric output, because LLMs are primarily designed for text-based tasks. However, the ground-truth thematic fit scores were averaged across the participants, providing the model with a graded scale, and therefore perhaps a numeric output would better correlate with our datasets. As hypothesized, our experiments using GPT  with categorical output yielded higher, although mostly modest increases in score – especially in Exp.2 and 4 on the Ferretti datasets (see Table 5 in the Appendix). The lower correlation scores for Llama  relative to GPT  could stem from different post-training methods (such as RLHF), but not enough information on this is publicly available. Therefore, we currently have no solid explanation for this difference between the models. See further analysis below in this section.
5.2 Analysis by Dataset Comparisons
Why were the tested LLMs much stronger on the Ferretti datasets than on Pado and McRae?
We see both models’ high performance on the Ferrerri datasets, and mediocre to low scores on Pado and McRae. Looking closely at 50 random samples from the GPT  output on Pado and McRae, we see and conclude the following:
GPT was weak at interpreting PropBank-based semantic roles: The semantic roles in Pado and McRae datasets are ProbBank semantic roles (Arg0, Arg1, Arg2). But the Ferretti datasets use the semantic roles Location and Instrument (Gruber, 1965; Fillmore, 1968), which are also commonly used words in English, unlike the PropBank notation. So either LLMs do not make the connection between Arg0, Arg1, Arg2 and their PropBank sense, or that LLMs are much weaker at PropBank roles, due to the distribution and much lower frequency in the training data. To test whether it’s because of the opaque PropBank sense, we conducted another experiment on the Pado and McRae datasets, adding to the prompt the word “PropBank” before the role, as follows:
“Given the predicate ’advise’, what properties should its PropBank Arg1  role have?”
This simple addition raised the models’ correlation scores on Pado from .25 to .28 (Exp.1.1 vs. Exp.3.1), and from .22 to .28 (Exp.1.2 vs. Exp.3.2). Similarly, correlation scores on McRae increased from experiments from .25 to .33 (Exp.1.1 vs. Exp.3.1), and from .27 to .36 (Exp.1.2 vs. Exp.3.2), supporting our opaque PropBank role hypothesis (while not ruling out other contributing factors).
Why did LLM performance differ greatly between Fer-Loc and Fer-Ins?
[…]
Why did generated sentences yield mixed results?
We observed that generated sentences with Step-by-Step Prompting lowered correlation with the Ferretti datasets but increased it on Pado and McRae, as shown in Table §2. By design, if the model determines a generated sentence incoherent, we back off to the corresponding lemma tuple input. Otherwise, it could hurt the thematic fit reasoning and conclusion in Exp.4.x vs. 3.x. We hypothesize that a large number of incoherent sentences slipped through and caused this drop on the Ferretti sets. Thus, we counted the slipped-through incoherent sentences in random samples of Fer-Loc and McRae, but found them to be 15% and 21%, respectively. Since these are similar ratios in the two tasks, this finding can not serve as an explanation. Still, this finding may guide future work in better generating and filtering sentences.
5.3 Analysis by Model
LLMs’ ability to "understand" natural language is likely enhanced by the grounding (of sorts) that pre-training code with comments provides. But in spite of Llama’s reported slight (.8%) advantage over GPT in code generation (Roziere et al., 2023), Llama  consistently performed worse than GPT  across all three axes of the experiments.
Why was Llama’s correlation on Pado and McRae near zero?
We don’t have a good explanation yet, and can only share general impressions: (1) Llama did not generate very diverse sentences. (2) Given tuples with the same predicate and arguments but different semantic roles, Llama did not distinguish between them with different scores. (3) Llama only provided "thin" reasoning, unlike GPT.
6 Conclusion and Future Work
We set out to discover if, or to what extent, LLMs such as GPT  and Llama possess linguistic and other knowledge required to estimate the thematic fit of event participants. We tested three axes for eliciting this knowledge: (1) Reasoning Form: adding Step-by-Step Prompting, (2) Input Form: adding generated sentences as context, and (3) Output Form: predefined categories vs. numeric scores.
Overall, using GPT  yielded higher results than using Llama, and new state-of-the-art scores were set on each test set. This indicates GPT acquired much of the linguistic knowledge required for thematic fit estimation. However, no single axis setting consistently yielded the highest scores. Step-by-Step Prompting helped much in both models on Fer-Loc, helped Llama but had mixed effect on Fer-Ins, and hurt GPT on Pado and McRae. Generating sentences helped GPT with Step-by-Step Prompting on Pado and McRae but had mixed effects elsewhere. Predefined categorical (compared to numeric) output raised GPT’s results across the board with few exceptions, but lowered Llama’s. In future work, we hope to improve Step-by-Step Prompting, following our analysis, especially by generating better sentences as context, and filtering out incoherent sentences more effectively. As bad generated sentences and bad reasoning greatly hurt results, we intend to explore using majority voting over many generated sentences, many filtering decisions, and many reasoning chains.
Future work should repeat our experiments, as new models are released with larger input windows, more parameters, and more effective post-training.
7 Limitations
Our experiments were evaluated on small datasets (a known limitation of this subfield), so findings and conclusions should be assessed cautiously. Our study considered only two LLMs, so one should be cautious about drawing conclusions for all LLMs. LLMs can be very sensitive to prompt phrasing, so it is possible that different prompts may yield very different results. The richness of prompt variation (prompt engineering) depends on human resources (different people use language differently) and funding/time resources, all of which were limited. Future work may achieve better results with more resources, if available. Due to the scarcity of evaluation data, we could not use prompt optimization methods such as DSPy Khattab et al. (2024).
Last, all findings and conclusions here rely on data and annotations in English, as it is by far the most resource-rich. But thematic fit (and semantics in general) are thought of as universal, and of course relevant to all languages. Still, we would advise to be cautious of drawing universal conclusions before validating our findings on additional, diverse set of languages.
Acknowledgments
The first author would also like to express gratitude to the Kuwait Chamber of Commerce and Industry for granting her a scholarship that made this research possible at Columbia University. The first author would also like to thank her family, especially her parents, for their support and motivation throughout this study. We are grateful to Google for providing us access to the Google Cloud Platform to conduct the experiments.
Appendices
Appendix A Prompt Design
Each experiment has a specific input, output, and reasoning design, as detailed in Sec. §3.1. The corresponding prompt design for each experiment is specified here.
A.1 Simple Prompting  with Head Lemma Tuples (Exp.1.1 — Exp.1.2)
[…]
  • Given the predicate ’eat’, how much does the argument ’pizza’ fit the role of Arg1?
As mentioned above, there are two output settings: numeric score (Exp.1.1) or categorical score (Exp.1.2). We extend the simple prompt as follows:
  • x.1: Numerical score where the model provides a numeric value between 0 and 1 to represent the thematic fit score:
    Reply only with a valid JSON object containing the key {"Score"} and a value that is a float number from 0 to 1, according to the following structure {{"Score"}: String}. Avoid adding any text outside this JSON object.
  • x.2: Categorical score where the model selects one of the predefined labels that indicate the thematic fit score:
    Reply only with a valid JSON object containing the key {"Score"} and a value that is one of ’Near-Perfect’, ’High’, ’Medium’, ’Low’ or ’Near-Impossible’, according to the following structure {{"Score"}: String}. Avoid adding any text outside this JSON object.
A.2 Simple Prompting  with Generated Sentences (Exp.2.1 — Exp.2.2)
The difference between experiments 1.x and 2.x is the input type. We augment the simple prompt with a natural language sentence previously generated by the model in the ‘sentence generation’ phase, from the same lemma tuples as in Exp.1.x:
  • Given the following sentence, ’I ate a pizza with my friends’, for the predicate ‘eat’, how much does the argument ‘pizza’ fit the role of Arg1?
A.3 Step-by-Step Prompting  with Lemma Tuples (Exp.3.1 — Exp.3.2)
Experiments 3.x are the enhanced version of the baseline Exp.1.x, where the reasoning steps precede the final semantic-fitness prompt. The reasoning phase is done through Step-by-Step Prompting, where a series of three prompts help reveal the reasons/internal knowledge that the model relies on to rate the thematic fit of the given data. These three prompts are:
  • Given the predicate ‘eat’, what properties should the its PropBank Arg1 role have? [^7] In an early version of this experiment, only the semantic role name was mentioned. After analyzing the early version as specified in Section §5, we suspected that the LLMs  may have not made the connection between roles such as ‘Arg1’ and their PropBank sense in this context. Therefore. we added the word ‘PropBank’ before the semantic role in the Pado and McRae datasets, as their semantic roles were ProbBank-based, unlike the Ferretti datasets. This indeed improved the results on these sets by 2-3% (absolute).
  • Given the predicate ‘eat’, and the argument ‘pizza’ in the Instrument role, what relevant properties does the argument have?
  • Given the above, how will the argument ‘pizza’ fit the PropBank Arg1  role for the predicate ‘eat’?
A.4 Step-by-Step Prompting  with Generated Sentences (Exp.4.1—Exp.4.2)
This last set of experiments combines Step-by-Step Prompting  in Exp.3.x and the generated sentence input type as in Exp.2.x. In the later version of this experiment [^8] In a preliminary version of experiment 4, the reasoning results of experiment 3 were reused. The results were .12-.15 higher on Fer-Loc, .04 higher on Fer-Ins, and .26-.33 lower on Pado and McRae., the first three prompts were as in Exp.3.x but with the addition of the generated sentence in each prompt. Then the simple thematic fit prompt follows, also with one of the generated sentences as context, as in Exp.2.x :
  • Given the following sentence, ‘I ate a pizza with my friends.’, for the predicate ‘eat’, how much does the argument ‘pizza’ fit the role of Arg1?
Appendix B Preliminary Study Results
Table 4 shows the results of the preliminary study that we conducted to choose the most suitable values for temperature and top_p of the models, following which we set temperature=0 and top_p=.95.
temp top_p .5 .7 .95
0 .77 .72 .80
.5 .52 .71 .75
.9 .61 .78 .74
Table 4: Preliminary Study for temperature and top_p; each cell represents the respective Spearman rank Correlation Coefficient ($\rho$)
Appendix C Output Form Result
Table 5 shows absolute improvements in GPT’s correlation scores when switching from numeric output to categorical output.
Exper. Dataset Fer-Loc Fer-Ins Pado McRae
Exp.1.1 vs. Exp.2.2 0.3 1.5 1.6 2.0
Exp.2.1 vs. Exp.2.2 8.8 4.6 1.4 2.8
Exp.3.1 vs. Exp.3.2 0.3 0.6 0.5 3.3
Exp.4.1 vs. Exp.4.2 4.3 2.2 1.0 2.2
Table 5: GPT improvement of categorical score over the numeric score. All values are absolute score differences.
Appendix D Analysis: Why did reasoning have mixed effects?
The achievements of Step-by-Step Prompting  in arithmetic, commonsense, and symbolic reasoning motivated us to test its potential in linguistic tasks. The linguistic tasks are similar to these arithmetic tasks in that they both require Step-by-Step Prompting, and having a (virtual or physical) scratchpad for these intermediate steps or calculations, helps get to the final answer. However, different tasks require different types of knowledge and reasoning capability. The thematic fit task requires a semantic reasoning ability and the ability to provide graded, non-binary judgment. Our experiments, conducted with this “scratchpad hypothesis” in mind, did not turn out as expected. As mentioned in Section §4, the effect of Step-by-Step Prompting  shines with the Fer-Loc dataset, but we observe the success of Simple Prompting  over Step-by-Step Prompting  on other datasets.
We analyzed 50 random results across all datasets from Exp.3 to address the question: Can good/bad reasoning (as defined below) positively/negatively affect the model’s performance on thematic fit estimation tasks?
We define the following four terms:
Good Reasoning: The type of reasoning that includes correct facts/properties about the specified role/argument and is logically valid with the given predicate, arguments, and arguments’ role.
Bad Reasoning: The type of reasoning that includes incorrect facts/properties about the specified role/argument and/or is logically invalid with the given predicate, argument, and argument’s role.
Good Scoring: The difference between the human raters’ (normalized) scores and the models’ scores is less than an arbitrary threshold.
Bad Scoring: The difference between the human raters’ scores and the models’ scores is greater than the threshold.
Classifying the reasoning result of each prompt in the sample as good or bad has been done manually. Classifying the thematic fit score result of each prompt in the sample as good or bad has been done according to the following steps:
  1. 1.
    Normalizing the human-based scores by converting them from a 1-7 scale to a 0-1 scale.
  2. 2.
    Comparing the difference between the normalized human-based scores to model-based scores for each test set item in the sample.
  3. 3.
    We experimented with multiple thresholds to classify the scores. This threshold is a heuristic calculated based on the six-interval normalized human scale. The first threshold is less than one interval (10%), the second is 1.5 the interval (25%), and the last one is more than three times the interval (50%). The most reasonable threshold was 25%, which is moderate between a strict threshold (10%) and a loose one (50%).
Reasoning Type Fer-Loc Fer-Inst Pado McRae
Score Type Num Cat Num Cat Num Cat Num Cat
Prompt 1 Good Reasoning 33:17 39:11 29:21 32:18 15:15 20:10 21:19 23:17
Bad Reasoning 0:0 0:0 0:0 0:0 8:12 8:12 8:2 8:2
Prompt 2 Good Reasoning 33:17 39:11 29:21 32:18 19:20 24:15 28:20 30:18
Bad Reasoning 0:0 0:0 0:0 0:0 4:6 4:6 1:1 1:1
Prompt 3 Good Reasoning 33:17 39:11 29:21 32:18 17:15 22:10 23:18 25:16
Bad Reasoning 0:0 0:0 0:0 0:0 6:12 6:12 6:3 6:3
Overall Good Reasoning 33:17 39:11 29:21 32:18 15:13 20:8 20:18 22:16
Bad Reasoning 0:0 0:0 0:0 0:0 8:14 8:14 9:3 9:3
Table 6: Good-to-bad score ratio across the board
’Num’ means numeric score, whereas ’Cat’ means categorical score.
The overall reasoning is considered ‘good’ if all prompts’ responses were ‘good’. Otherwise, it is considered bad reasoning.
The result of this analysis is shown in Table 6. On Pado and McRae, most cases have a high mass of ‘good scores’ as opposed to ‘bad scores’ with good reasoning. Cases of bad reasoning were fewer and far apart. Also, the good-to-bad score ratio in good reasoning was much higher compared to that ratio in bad reasoning. On the Ferretti sets, there were no bad reasoning cases, but the same good-to-bad score ratio pattern was observed. This pattern is reflected in both numeric and categorical scores, with minor differences.
Returning to our question: Can good/bad reasoning (as defined above) positively/negatively affect the model’s performance on thematic fit estimation tasks?
[…]
Figure 3: Effect of Early Bad Reasoning
Appendix E Analysis: Why did LLM performance differ greatly between Fer-Loc and Fer-Ins?
To answer this question, we analyze 50 random samples from the GPT  output on Fer-Loc and Fer-Ins.
[…]
Prompt-1 (Instrument  properties) Response:
For the predicate ’float’, the Instrument  role should have properties indicating that it is an object or mechanism that enables or assists the action of floating. This could include buoyancy aids, flotation devices, or any tool that helps an entity to remain on or move through the surface of a liquid without sinking.
Prompt-2 (wood properties) Response:
The argument ’wood’ has properties of buoyancy and low density relative to water, which allow it to remain on or move through the surface of a liquid without sinking.
Prompt-3 (thematic fit) Response:
The argument ’wood’ fits the Instrument  role for the predicate ’float’ well, as its inherent buoyancy and low density are properties that enable or assist in the action of floating.
This indicates that the model has acquired sufficient linguistic and world knowledge to reliably estimate thematic fit for the Instrument  role, at least for some cases.
[…]
Prompt-1 (Location  properties) Response:
The Location role for the predicate ’study’ should refer to a place where the studying activity occurs. This could include physical locations such as a library, a classroom, a home office, or a coffee shop. It could also encompass more abstract locations in a digital sense, such as an online learning platform or a virtual classroom. The key properties are that the location supports, or is conducive to, the activity of studying, providing an environment where learning materials can be accessed, and concentration is facilitated.
Prompt2 (‘mall’ properties) Response:
The argument ’mall’ for the predicate ’study’ is unconventional as malls are typically associated with shopping, dining, and entertainment rather than studying. However, relevant properties could include the presence of seating areas, food courts, or coffee shops within the mall where studying could potentially occur. Additionally, malls are generally public, accessible spaces that might offer a change of scenery for someone looking to study outside of more traditional environments like libraries or home offices.
Prompt3 (thematic fit) Response:
The argument ’mall’ fits the Location role for the predicate ’study’ in a non-traditional or unconventional manner. While not primarily designed for studying, a mall can offer spaces such as seating areas or cafes that could be repurposed for studying. The fit is atypical and might not be ideal due to potential distractions, but it is possible under certain conditions.
This indicates that the model has acquired sufficient knowledge to reliably estimate thematic fit for the Location role, at least for some cases, even with some unconventional cases (here, ‘mall’ for studying).
[…]
Prompt-1 (Instrument  properties) Response:
The Instrument role for the predicate ’cut’ should have properties indicating that it is capable of dividing or separating an object into parts or portions, typically through a sharp edge or by applying force. It should suggest an object or tool designed or used for cutting, such as a knife, scissors, or saw.
Prompt-2 (‘razor’ properties) Response:
The argument ’razor’ has properties indicating it is a tool designed for cutting, specifically for shaving hair from the skin. It has a sharp edge capable of making precise cuts, making it suitable for the Instrument role in the context of the predicate ’cut’.
Prompt-3 (thematic fit) Response: The argument ’razor’ fits the Instrument role for the predicate ’cut’ very well, as it possesses the necessary properties of having a sharp edge designed for precise cutting, aligning with the requirements of an instrument used for cutting.
This indicates that evaluating thematic fit for the Instrument role may require more complex reasoning. In this example, a razor is a candidate instrument for cutting, but a razor is a specialized instrument for cutting beards or other kinds of body hair. People typically do not use ‘cut’ for this action, even if they describe shaving the head. Moreover, people typically do not use a razor to cut fruit (even more so with modern shaving sticks, where the razor is embedded in the stick or the add-on cartridge). Thus, more complex reasoning is needed to discover other related properties required. The model’s thematic fit score here was high, but it should have been in the mid-range.
The Fer-Loc dataset is more affected by Step-by-Step Prompting  than the Fer-Ins. When we measure the score difference between Simple Prompting  and Step-by-Step Prompting  (Exp.3.1 vs. Exp.1.1), 68 tuples (28%) in Fer-Ins had no difference. However, in Fer-Loc, only 53 tuples (19%) showed no difference in scores.
Our findings show that reasoning about the Location  is often much simpler than reasoning about the Instrument, which may also require complex reasoning about the compatibility of the physical properties of the Instrument  to the Patient or Theme given the predicate, and/or the linguistic definition or specialization (shaving but not cutting, as in example c above). Even though in principle this could be needed for reasoning about Location too, it seems to be less common in practice. Thus, we conclude that the tested model can reason about the thematic fit for Location  much better than for an Instrument, because the model lacks the additional capabilities or knowledge required for the latter role (at least without more specialized prompting). This also explains the improvement from Exp.1 to Exp.3 in Fer-Loc.
Appendix F Details of Models and Hardware
Closed model
We chose OpenAI GPT-4-turbo (gpt-4-0125-preview) (Achiam et al., 2023) due to the GPT family’s outstanding performance history in downstream NLP tasks. GPT models have achieved high performance since the first version, GPT-1, achieving state-of-the-art results in 9 out of 12 NLP tasks (Radford et al., 2018). This still holds for the latest versions (Bubeck et al., 2023). GPT-4-turbo has a 128,000 token input buffer and was trained on data up to April 2023. We used the OpenAI API to access the model on two vCPUs with 7.5 GB memory on the Google Cloud Platform (GCP) (https://cloud.google.com/gcp).
Open model
We chose Meta CodeLlama2 (codeLlama2-70B-instruct) (Roziere et al., 2023) with 100,000 token input buffer because it is a high-performing open-source model for prompting. Adapting rotary positional embedding (RoPE) in CodeLlama enhances the model’s understanding of the context (Roziere et al., 2023). Moreover, it showed competitive results compared to GPT-3.5, including its zero-shot learning ability (Roziere et al., 2023). Last, models pretrained on code have shown better performance on semantic tasks (Petty et al., 2024). CodeLLama2 is trained on proprietary non-public instruction data. As the number of parameters is vast (13B), it requires 13 billion x bfloat16 = 26GB of GPU memory. Latency is another major challenge as the prompt series increases in length. Thus, to achieve good inference time, we use VLLM (Kwon et al., 2023) with PagedAttention, on a GCP virtual machine with 24 vCPUs with 96GB memory, and 2 GPUs with 48GB GDDR6.  [^10] These models were state-of-the-art and the latest available to us at the time of experimentation.