### v2

# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

Abstract
We show closed models possess much thematic fit knowledge and set a new state of the art, while open models also seem to capture much relevant knowledge (in semantic filtering), but yield lower scores. Surprisingly, multi-step reasoning only helped closed models (with few exceptions); generated sentences hurt closed models’ performance; and output form had little to no effect. We analyze the reasons for these findings, and conclude that more foundational work is needed for a single LLM to perform the best on all tasks with the same experimental condition, let alone improve results further. Source code is available at: https://github.com/SafeyahShemali/LLM_Thematic_Fit_25
Uncovering Autoregressive LLM Knowledge of Thematic Fit
in Event Representation
Safeyah Khaled Alshemali Imperial College London London, UK ska24@ic.ac.uk          Daniel Bauer Columbia University New York, NY, USA db2711@columbia.edu          Yuval Marton University of Washington WA, USA ymarton@uw.edu
1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge (including linguistic knowledge) of event arguments’ thematic fit to their assigned semantic role for a given predicate? A naive approach of prompting LLMs might yield negative or weak results. However, we hypothesize that this knowledge does exist, at least to a certain degree, but needs eliciting. We test our hypothesis along three axes: (1) Reasoning: Does structured reasoning in prompting improve the model’s estimations? [^1] Following chain-of-thought (wei2022chain). (2) Input: Does providing generated sentences as contexts help, compared to providing only a triplet of keywords? [^2] Traditionally consisting of the predicate’s lemma, the argument’s syntactic head’s lemma, and its semantic role. (3) Output: Does specifying categorical output (such as “low” or “high”) yield better results than a numerical output $(0..1)$ as the model’s thematic fit estimation?
Thematic Fit
The thematic fit estimation task measures the compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a specific semantic role assigned to the argument. [^3] Thematic fit is related to, but separate from, semantic role labeling (SRL). Given a sentence, SRL should output the semantic frames in the sentence, including the predicates, arguments, and their semantic roles in the frame. Thematic fit takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For theoretical and historical details, see gruber1965studies; fillmore1968lexical; dowty1991thematic; parsons:1990; mcrae1998modeling inter alia. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (usually denoted as Arg0 in PropBank; see sec.§3.2), then ‘people’, ‘cat’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1 in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. And so on. Psycholinguistic experiments show that humans rely on generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (mcrae2009people; bicknell2010effects).
Thematic fit estimation applies data-driven event representations and computational methods to mimic, or align with this type of human knowledge of events. This is challenging due to the lack of direct labeled training data. Thematic fit norms (see §3.2), which summarize human judgments (aka ratings), provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, thematic fit models are traditionally trained on proxy data (either PropBank (kingsbury2002treebank) or FrameNet (johnson2016framenet) annotations). Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (levy2017semantic), similarity judgments (hill2015simlex), and more. However, on thematic fit, these models, including masked LMs such as BERT (devlin2019bert), have so far only yielded modest improvements over previous approaches (pedinotti2021cat). In this work, we examine if more recent, large autoregressive models, such as GPT (classicalGPT; radford2018improving; achiam2023gpt), can produce better results via prompting. We attempt to elicit their relevant linguistic knowledge along the following three axes:
Axis 1: Reasoning (Simple vs. Step-by-Step Prompting)
Several studies show LLMs perform remarkably well in challenging NLP tasks such as Planning (wang2022towards), automatic scoring (lee2024applying), and question answering (wang2023keqing), when using ‘chain-of-thought prompting’ (wei2022chain). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise. [^4] It proved so influential that several leading labs have developed ‘thinking’ models with ’test-time compute’, incorporating this approach in the post-training and settings. We view thematic fit estimation as a linguistic reasoning task because it can be broken into clear sub-steps (see §3.1). Thus, we compare the LLMs’ performance using a “naive” single-prompt approach (Simple Prompting hereafter) with the chain-of-thought approach (Step-by-Step Prompting hereafter) on the thematic fit estimation task.
Axis 2: Input (Generated Sentences vs. Lemma Tuples)
[…]
Axis 3: Output (Predefined Categories vs. Numeric Score)
The thematic fit ratings in our test sets are Likert scale ratings (following mcrae1998modeling) averaged over 7 human raters. Thus, a straightforward evaluation would prompt the model to output numeric scores as well. However, LLMs were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. Moreover, the LLMs’ numeric scores are actually a textual output, which we then convert to its corresponding numeric value. Therefore, we also prompt the LLMs to rate the thematic fit according to predefined categories (“Low”, “High”, etc.), which we then convert to corresponding predefined numeric values for evaluation purposes (see Section §3.2). [^6] We leave for future work whether it is useful to use LLMs’ logprobs for this task, and how to best do that.
Contributions
Our main contributions are:
  • First study to prompt autoregressive LLMs in order to directly estimate thematic fit.
  • New state-of-the-art thematic fit estimation.
  • Comparison of three axes: reasoning form, input form, and output form.
  • A novel semantic filtering process to exclude semantically incoherent or semantically incompatible generated sentences.
  • Showing and analyzing the weakness of new vs. older GPT models in this task, and why chain-of-thought is not always helpful here.
2 Related Work
Early distributional approaches to the thematic fit task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (erk2010flexible; baroni2010distributional; santus2017; sayeed2015exploration; greenberg2015improving). Our first baseline (hereafter BG) is the distributional approach by greenberg2015improving, which is based on the syntactically structured distributional semantic model by baroni2010distributional. More recent approaches use neural networks to learn event representations. tilk2016event describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. hong2018learning improve over tilk2016event by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1). Marton2021ThematicFB replicate hong2018learning’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). muthupari2022s focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the hong2018learning model with random vectors vs. pre-trained GloVe embeddings (pennington2014glove), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
Contemporary work using pretrained LLMs (without any additional supervised training) for thematic fit either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’) (kauf2023event), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence metheniti2020relevant; pedinotti2021cat. Notably, pedinotti2021cat only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (chersoni2019structured). Finally, testa2023we examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis.
We are the first to use prompting with autoregressive LLMs directly for thematic fit.
Figure 1: Experiment Method. For more details of the Step-by-Step Prompting and Semantic Filtering, see §3.1. The output of Exp.3.x - Exp.4.x contains the score’s justification in addition to the score.
Reasoning Form Input Form Output Form Exp
Simple Prompting Head Lemma Tuples Num. Score 1.1
Cat. Score 1.2
Generated Sentences Num. Score 2.1
Cat. Score 2.2
Step-by-Step Prompting Head Lemma Tuples Num. Score 3.1
Cat. Score 3.2
Generated Sentences Num. Score 4.1
Cat. Score 4.2
Table 1: Overview of experiments. ’Num.’ = numeric score. ’Cat.’ = categorical score.
3 Experiments
The experiments were designed to investigate three axes (§1) with two possible values each. Therefore, we conduct $2^{3}=8$ experiments. We refer to them as ’Exp $n.m$’ (see Table 1). Each experiment has a prompting template according to the reasoning, input, and output settings, as detailed in Appendix §A. Consistent with the order in the introduction, we will next detail the design for the reasoning form, the input form, and the output form.
3.1 Method: Reasoning, Input and Output
Axis 1: Reasoning (Simple Prompting vs. Step-by-Step Prompting)
The first axis is concerned with the effectiveness of Step-by-Step Prompting (Exp.3.x and 4.x) compared to Simple Prompting (Exp.1.x and 2.x) in evaluating the thematic fit (leftmost column in Table 1). In the Simple Prompting approach, the model is asked directly only one prompt, which we call a ‘simple thematic fit prompt’, to score the thematic fit of some predicate’s argument to its assigned semantic role (where the predicate, the argument, and the semantic role are all given).
[…]
Axis 2: Input (Lemma Tuples vs. Generated Sentences)
[…]
The sentence must pass all the steps of the filtering process in at least one trial, or it will be filtered out. For each of the five sentences, the final fit score is averaged. If steps (1) or (2) fail, we back off to using the lemma tuples. The second column in Table 1 specifies the form of the input.
Axis 3: Output (Categorical vs. Numeric Scoring)
Each experiment may generate one of two forms of output: a numeric thematic fit score (Exp.x.1 in Table 1) or a predefined thematic fit category (Exp.x.2). The numeric thematic fit score (numeric score, hereafter) is a numeric value in the range [0,1]. The predefined thematic fit categories (categorical score, hereafter) are ’Unlikely/Impossible’, ’Low’, ’Medium’, ’High’, or ’Perfect’. These output score categories are then converted to the following numeric values: (0.0, 0.25, 0.5, 0.75, 1.0), respectively, in a post-processing script, for evaluation purposes (see §3.2). In Exp.3.x and Exp.4.x, the model is asked to generate a justification for the score it assigns to the input, based on its previous Step-by-Step Prompting reasoning output.
3.2 Datasets and Evaluation
Training set:
We used pre-trained LLMs (see sec. §3.4) “off the shelf” without modification. In an ideal world, we would be able to point the reader to the full pre-training data. We did not add psycholinguistic (or other) labeled data to the training set, nor did we fine-tune the LLMs.
Test sets:
[…]
mcrae1998modeling, hereafter ’McRae’, contains $1,444$ predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0 and Arg1 (kingsbury2002treebank). We removed eight duplicate records.
pado2006modelling, hereafter ’Pado’, contains 414 tuples in the same form as mcrae1998modeling; however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred native English speaker volunteers judged tuples in four randomly ordered lists.
ferretti2001integrating evaluate thematic fit for the roles Instrument (hereafter, the Fer-Ins dataset) and Location (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in mcrae1998modeling. However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
Evaluation:
For each test set, we used Spearman’s Rank Correlation Coefficient ($\rho$) (spearman:1904) between two lists: the test set, sorted by the average human ratings for each item, and the corresponding model output (sorted as well). Categorical output was converted to numeric values, so it can be sorted similarly too before conducting this evaluation (see §3.1).
3.3 Hyperparametes
We conducted a preliminary study to determine good temperature and top_p parameter values for the LLMs. We used GPT-3.5 on a subset of the Fer-Ins dataset (20 samples $\sim$ 8%) and ran a parameter sweep over temperature values (0.0, 0.5, 0.9) and top_p values (0.5, 0.7, 0.95). Temperature values were chosen near-uniformly across the scale to examine a wide range of the model creativity setting. Higher values correspond to higher randomness (aka "creativity") in the output. High values of top_p limit the randomness in the output to more probable tokens. Since we were more interested in accurate output, and less interested in creative writing for our purposes here, we tested top-p threshold values on the higher half of the scale. Temperature 0.0 and top_p 0.95 showed the highest correlation coefficient with human judgments on our sample. We set the max_tokens parameter based on the length of the prompt and its expected output. For experiments with Lemma Tuples input, we set max_tokens to 100, whereas for Generated Sentences input, we set it to 300 (for Simple prompting). As Step-by-Step Prompting is expected to be lengthy to clarify the underlying reasons, we set max_tokens to 600 there. We used these settings in all experiments.
3.4 Models and Hardware
We conducted the experiments using closed and open models, the latest available at the time of running the experiments. Closed models (as they are often best performing, but less reproducible): OpenAI gpt-4-0125-preview (achiam2023gpt; openai2023gpt4turbo) with unconfirmed 1.76 trillion parameters (hereafter ‘GPT4-Turbo’) and gpt-4.1-2025-04-14 (openai2025gpt41docs; OpenAI_GPT-4.1_2025) with 1.8 trillion parameter (hereafter ‘GPT4.1’). Open models (more controlled and reproducible, even if not fully open): llama3.2:3b-instruct-q4_K_M (Dubey2024Llama3; meta2024llama), a quantized version of Meta Llama3.2 with 3 billion parameter (hereafter ‘Llama’) and qwen2.5:7b-instruct-q4_K_M (Yang2024Qwen2.5) with 7.61 billion parameters (hereafter ‘Qwen’). Closed model inference utilized the OpenAI API. Open model inference utilized Ollama (Ollama2023), as it is resource-efficient.
4 Results
Overall, our results set new state of the art for all four tasks, especially with GPT4-Turbo using Step-by-Step Prompting (except for Pado, where GPT4-Turbo tied with the highest baseline). See Table 2 and details below.
4.1 Closed Models
Both closed models (GPT4-Turbo and GPT4.1), when using Step-by-Step Prompting, outperformed the baselines by a large margin, except for Pado (where GPT4-Turbo tied with the highest baseline, and GPT4.1 was weaker than all baselines). With respect to the three axes, the models’ performance differed as follows:
Axis 1: Reasoning Form
GPT4-Turbo demonstrated superior performance across the board and among all models, including GPT4.1, see Table 3. Step-by-Step Prompting experiments (Exp.3.x and 4.x) yielded the best scores, where absolute improvement over the strongest baseline was .23 for Fer-Loc, .22 for Fer-Ins, .23 for McRae, and a tie with Pado. Simple Prompting also outdid the baselines (except Pado).
GPT4.1 Step-by-Step Prompting also outdid all baselines except Pado, but curiously lower than GPT4-Turbo, in spite of being newer. Simple Prompting gave mixed results w.r.t. the baselines.
Axis 2: Input Form
Using Generated Sentences did worse or similar to using Lemma Tuples in all tasks and LLMs, except Pado with GPT4-Turbo.
Axis 3: Output Form
Using predefined categories compared to numeric output (Exp.x.2 vs. Exp.x.1) had zero, or small mixed effect, contrary to our hypothesis.
With GPT4-Turbo, using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) leads to a mixed effect across the board. Regardless of input and reasoning forms, the categorical score is lower than the numerical score in Fer-Loc. For Fer-Ins and Pado, the categorical score is either the same or higher than the numerical score, as the gain in Fer-Ins was .04 in Exp2 and in Pado was .04 in Exp3. Using categorical output in McRae lowers the score in general, except for Exp4, where the gain was .02.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .63 .64 .35 .54
Exp.1.2 .61 .65 .40 .58
Exp.2.1 .48 .49 .42 .55
Exp.2.2 .46 .54 .45 .56
Exp.3.1 .69 .70 .48 .65
Exp.3.2 .68 .68 .46 .66
Exp.4.1 .57 .66 .58 .58
Exp.4.2 .58 .66 .59 .59
Table 2: GPT4-Turbo Spearman Rank Correlation ($\rho$). BG = greenberg2015improving, B0 = tilk2016event, B1 = hong2018learning, B2 = Marton2021ThematicFB, B3 = muthupari2022s as baselines for our experiments. Subscripts a-c denote specific models there (Glove-shared-not-tuned, etc.). Exp.m.nm.n = see Table 1. All $\rho$’s had p-values $10^{-13}$.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .51 .56 .40 .64
Exp.1.2 .50 .56 .40 .58
Exp.2.1 .32 .33 .33 .40
Exp.2.2 .26 .37 .33 .37
Exp.3.1 .59 .59 .45 .64
Exp.3.2 .56 .59 .42 .63
Exp.4.1 .56 .53 .42 .44
Exp.4.2 .53 .53 .42 .46
Table 3: GPT4.1 Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-5}$.
4.2 Open Models
The performance of the open models was lower than the closed models, and lower than several baselines across the board, except for Fer-Loc using Step-by-Step Prompting with Llama and McRae with Qwen (mostly in Exp.4.x. Details are below:
Axis 1: Reasoning Form
Despite being mostly lower than the baselines, Step-by-Step Prompting has a positive effect in Fer-Loc and McRae. In Pado, the effect was detrimental with Llama (Table 4), but positive with Qwen (Table 5).
Axis 2: Input Form
Using Generated Sentences instead of Lemma Tuples had mostly no effect with Llama, except for gains in McRae. With Qwen there were mostly gains except for lower scores in Exp.4.x in FerLoc.
Axis 3: Output Form
Using predefined categories in the output, compared to numeric output (Exp.x.2 vs. x.1), yielded lower scores with Llama, except in Pado Exp.3.x and 4.x, but the correlation scores there are too low to trust these differences are meaningful. With Qwen  results were worse in Fer-Ins and McRae, and mixed elsewhere.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .32 .33 .27 .16
Exp.1.2 .30 .23 .26 .13
Exp.2.1 .32 .33 .31 .25
Exp.2.2 .29 .22 .22 .23
Exp.3.1 .46 .29 .05* .23
Exp.3.2 .41 .24 .10* .20
Exp.4.1 .46 .28 .06* .27
Exp.4.2 .41 .24 .12* .21
Table 4: Llama Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-3}$ except cells with * for which p-values were between (.01 - .22).
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .29 .43 .28 .42
Exp.1.2 .38 .37 .19 .34
Exp.2.1 .31 .44 .31 .48
Exp.2.2 .37 .38 .22 .39
Exp.3.1 .41 .36 .32 .40
Exp.3.2 .41 .32 .36 .35
Exp.4.1 .31 .37 .33 .51
Exp.4.2 .22 .36 .37 .47
Table 5: Qwen Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-3}$.
5 Discussion
The core questions we aim to address in this work are: Do LLMs have the kind of linguistic knowledge required for thematic fit estimation? If so, (Q1) does Step-by-Step Prompting help utilize the model’s internal linguistics knowledge for this task? (Q2) Would providing more context by including sentences as part of the input help? (as opposed to providing only lemma tuples with semantic roles) (Q3) Do predefined thematic fit output categories better elicit the model’s internal linguistic knowledge compared to numeric output?
Overall, we conclude that closed models acquired much of the linguistic knowledge required for thematic fit estimation. Results with both GPT models on each of the test sets achieved or even surpassed all previous baselines, which were trained on labeled linguistic data (even if not directly trained for thematic fit estimation). This is impressive, [^9] The available inter-annotator agreement or correlation information for our testsets indicates the upper bound for these tasks is between the mid-60’s and mid-70’s. GPT4-Turbo’s scores got close, but no single setting was always the best. assuming closed models were not trained on this task (sadly, we cannot know for sure because they are, well, closed). Next, we discuss the findings by the three axes (as defined in §3) and then move to analyzing additional aspects.
(Q1) Axis 1: Reasoning Form
For GPT models (closed models), Step-by-Step Prompting fulfilled its intended goal of providing a “scratchpad” to help them calculate inferences given their internal knowledge (with the exception of GPT4.1 in Pado). In contrast, Step-by-Step Prompting didn’t help with open models (except for Fer-Loc with Llama and for McRae with Qwen).
Why was there a difference in the Step-by-Step Prompting effect between the closed and open models? Step-by-Step Prompting is designed to help the models reason by first eliciting useful properties of arguments and roles, before the final thematic fit estimation. However, our analysis on a sample of random tuples revealed that open models are weak at inferring the essential properties that are most expected for an argument, to serve well in a specific semantic role. Input from this step, containing invalid or irrelevant properties, often derailed the final reasoning output quality (See Appendix §E).
Why did GPT4-Turbo outperform GPT4.1 in spite of being older? We observed that GPT4.1’s ratings were often more ‘cautious’, meaning it rarely output very high or very low scores. We verified this by calculating the variance of each model’s output for each task and experimental setting (Appendix §B, Tables 7-8), finding a negative correlation between the variance and the correlation score ($\rho$). The lower variance of GPT4.1 means it is more likely to do worse on clear-cut inputs (with very high or very low thematic fit).
(Q2) Axis 2: Input Form
In principle, providing more context enhances models’ ability to comprehend complex tasks. This was generally reflected in our experiments with open models, but not closed models (with few exceptions).
Why do generated sentences only help open models? Although the open models did not outperform the closed models, they often better leveraged the input with generated sentences, and improved the correlation score ($\rho$) relative to the tuple input. But sentence input is likely to outperform tuple input only if these sentences are “good” (semantically coherent, with the given argument in the given role for the given predicate). Moreover, “bad” sentences are likely to derail the models’ performance. Therefore, we analyzed a random sample of 20 tuples, 5 generated sentences for each, altogether 100 sentences per model. We found that the open Qwen did better in filtering out “bad” sentences (in spite of having fewer parameters): its semantic filtering let through no False Positives (FPs), thus successfully leveraging the sentence inputs. In contrast, the closed GPT4-Turbo let through 6 FPs (Table 6). We saw similar results comparing filtering by Qwen vs. GPT4-Turbo, both on GPT4-Turbo’s sentences. [^10] Note that Qwen had higher False Negative rate, but this would only result in regressing to tuple input (§3.1 Axis 2). With it, closed models clearly generated more “good” sentences (80 vs. 60 True Positives + False Negatives in our sample). Therefore we caution against jumping into conclusions which model gained more relevant knowledge.
Acceptable Sentence? Qwen GPT4- Turbo Qwen filtering GPT4-Turbo
True Positive 18 52 26
True Negative 40 15 20
False Positive 0 6 1
False Negative 42 27 53
Table 6: Semantic filtering of “bad” generated sentences, random sample size: 100. “Bad” = True Negatives + False Positives. Left 2 columns: model both generated and filtered.
(Q3) Axis 3: Output Form
There are arguments in favor of either categorical and numeric output. On the one hand, we expect predefined output categories to yield better results than numeric output, because LLMs are primarily designed for text-based tasks. On the other hand, in spite of this, LLMs have also shown impressive performance in math (even if not perfect). Also, the ground-truth thematic fit scores were averaged across the participants, providing a graded scale, and therefore perhaps a numeric output would better correlate with that. Indeed, output form had mostly small-to-zero effect (up to 0.02 difference) for closed models, but categorical output gave mixed results, mostly worse than numeric output in open models. We hypothesize this may be due to the differences in post-training between the closed and open models, but we currently have no solid explanation for this difference, and we leave this for future work.
6 Conclusion and Future Work
We set out to discover if, or to what extent, autoregressive LLMs possess linguistic and other knowledge required to estimate the thematic fit of event participants. The tested LLMs set new state-of-the-art on each test set (and tied on Pado). Still, our takeaway is that thematic fit is not yet “solved”, in spite of this clear progress. Closed models [^11] We observed that GPT4.1 did worse than GPT4-Turbo, in spite of being newer, likely because 4.1 is too ‘cautious’ (§5). yielded higher results than open models, indicating that closed models, especially GPT4-Turbo, acquired much of the linguistic knowledge required for thematic fit estimation. However, no single axis setting consistently yielded the highest scores. Moreover, open models where better at filtering “bad” generated sentences, so we caution against rushing to interpret the results as showing that open models gained less relevant knowledge. It is still unclear if models need additional knowledge, or if the right elicitation has not been found yet.
We tested three axes for eliciting this knowledge: (1) Reasoning Form: adding Step-by-Step Prompting, (2) Input Form: adding generated sentences as context, and (3) Output Form: predefined categories vs. numeric scores.
(1) Surprisingly, Step-by-Step Prompting only helped closed models; Our analysis showed that invalid or irrelevant reasoning not only was unhelpful but actually hurt performance. This happened more in open models.
(2) Generated sentences hurt closed models; We saw LLMs struggle with generating sentences where the given argument is a better fit to a semantic role different than the given one. To tackle this, we used novel semantic filtering, which, to our surprise, worked better for open models, because the latter were better at filtering out False Positives.
(3) Predefined categorical output had a mixed effect (relative to numeric output): zero-to-small effect with closed models and more negative effect with open models.
In future work, we hope to improve Step-by-Step Prompting, following our analysis, especially by generating better sentences as context, and better filtering out “bad” sentences. If successful, LLMs may also be used to study issues in thematic fit, in line with futrell2025how, BabyLM efforts [^12] https://babylm.github.io, and others who believe that LLMs can be useful tools to study linguistic phenomena.
Limitations
Our experiments were evaluated on small datasets (a known limitation of this computational psycholinguistics subfield, as opposed to typical machine learning), so findings and conclusions should be assessed with caution. Our study considered only few LLMs, so one should be cautious about drawing conclusions for all LLMs.
LLMs can be very sensitive to prompt phrasing, so it is possible that different prompts may yield very different results. Also, prompt engineering is tricky; thus, tweaking the prompts may lead to different results. The richness of prompt variation (prompt engineering) depends on human resources (different people use language differently) and funding/time resources, all of which were limited. Future work may achieve better results with more resources, if available. Due to the scarcity of evaluation data, we could not use prompt optimization methods such as DSPy Khattab2023DSPyCD or fine-tune the models.
Last, all findings and conclusions here rely on data and annotations in English, as it is by far the most resource-rich. But thematic fit (and semantics in general) are thought of as universal, and of course relevant to all languages. Still, we would advise to be cautious of drawing universal conclusions before validating our findings on an additional, diverse set of languages.
Appendices
Appendix A Prompt Design
Each experiment has a specific input, output, and reasoning design, as detailed in Sec. §3.1. The corresponding prompt design for each experiment is specified here.
A.1 Simple Prompting with Head Lemma Tuples (Exp.1.1 — Exp.1.2)
[…]
  • Given the predicate ‘eat’, how well does the argument ‘pizza’ fit the PropBank Arg1 role?
As mentioned above, there are two output settings: numeric score (Exp.1.1) or categorical score (Exp.1.2). We extend the simple prompt as follows:
  • x.1: Numerical score where the model provides a numeric value between 0 and 1 to represent the thematic fit score:
    Reply only with a valid JSON dictionary {"Score": numeric_value} where numeric_value is a float number from 0 to 1. Avoid adding any text outside this JSON dictionary.
  • x.2: Categorical score where the model selects one of the predefined labels that indicate the thematic fit score:
    Reply only with a valid JSON dictionary {"Score": string_value} where string_value that is one of "Perfect", "High", "Medium", "Low" or "Unlikely/Impossible". Avoid adding any text outside this JSON dictionary.
A.2 Semantic Filtering Process for Generated Sentences
Before executing experiments 2 and 4, the model is asked to generate $k$ diverse semantically coherent sentences containing an argument in a specific semantic role, given a predicate. Here we use $k=5$.
  • Give me five sentences with the predicate ‘eat’ and argument ‘pizza’ in the PropBank Arg1 role. Make the sentences as diverse as possible, while using the predicate as the main verb and keeping the given argument in the specified role. Make sure that each sentence is semantically coherent and the argument is in the given role.
LLMs may often generate a semantically incoherent or incompatible sentence with respect to the given role for the argument. Thus, we design a semantic filtering process of 4 waterfall Yes/No Steps. If the sentence passed the whole waterfall at least once out of three trials, it is considered semantically acceptable. Here are the prompts we used:
  • (2a) Here’s a sentence: ‘I ate a pizza with my friends’. Is the given sentence semantically coherent?
  • (2b) For the sentence: ‘I ate a pizza with my friends’. Does the sentence contain the predicate ‘eat’?
  • (2c) For the sentence: ‘I ate a pizza with my friends’. Does the sentence contain ‘pizza’ as an argument of the predicate ‘eat’?
  • (2d) For the same sentence: ‘I ate a pizza with my friends’. Given the definitions, instructions, and answers above, is ‘pizza’ in the PropBank Arg1 role for the predicate ‘eat’?
A.3 Simple Prompting with Generated Sentences (Exp.2.1 — Exp.2.2)
The difference between experiments 1.x and 2.x is the input type. We augment the simple prompt with a natural language sentence previously generated by the model in the ‘sentence generation’ phase, from the same lemma tuples as in Exp.1.x:
  • Given the following sentence, ‘I ate a pizza with my friends’, how well does the argument ‘pizza’ of the predicate ‘eat’ fit the PropBank Arg1 role?
A.4 Step-by-Step Prompting with Lemma Tuples (Exp.3.1 — Exp.3.2)
Experiments 3.x are the enhanced version of the baseline Exp.1.x, where the reasoning steps precede the final semantic-fitness prompt. The reasoning phase is done through Step-by-Step Prompting, where a series of four prompts helps reveal the reasons/internal knowledge that the model relies on to rate the thematic fit of the given data. These four prompts are:
  • What salient essential properties or core characteristics best describe the essence of ‘pizza’? [^13] In an early version of this experiment, only the semantic role name was mentioned. After analyzing the early version, we suspected that the LLMs may have not made the connection between roles such as ‘Arg1’ and their PropBank sense in this context. Therefore. We added the word ‘PropBank’ before the semantic role in the Pado and McRae datasets, as their semantic roles were PropBank-based, unlike the Ferretti datasets. This indeed improved the results on these sets by 2-3% (absolute).
  • Given the predicate ‘eat’, and the argument of it in the PropBank Arg1 role, what salient essential properties or core characteristics are most expected for this argument, to serve well in the PropBank Arg1 role?
  • Given the predicate ‘eat’, what PropBank roles are most suitable or likely for its argument ‘pizza’?
  • Given the predicate ‘eat’ with the argument ‘pizza’ in the PropBank Arg1 role, what essential properties does the argument have, properties which fit well its PropBank Arg1 role? And what essential properties this argument should have had (but are missing) in order for it to serve well in its PropBank Arg1 role?
A.5 Step-by-Step Prompting with Generated Sentences (Exp.4.1—Exp.4.2)
This last set of experiments combines Step-by-Step Prompting in Exp.3.x and the generated sentence input type as in Exp.2.x. In the later version of this experiment, the first four prompts were as in Exp.3.x but with the addition of the generated sentence in each prompt. Then the simple thematic fit prompt follows, also with one of the generated sentences as context, as in Exp.2.x :
  • Here’s a sentence: ‘I ate a pizza with my friends.’, What salient essential properties or core characteristics best describe the essence of ‘pizza’?
Appendix B Experimental Results with Variance
Table 7 and 8 show the Spearman correlation scores along with the variance for each experiment across the board. These tables support our explanation for the difference in performance between GPT4-Turbo and GPT4.1.
Model Dataset Fer-Loc Fer-Ins Pado McRae
$\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$
BG: GSD2015 .29 - .42 - .53 - .36 -
B0: NN RF .44 - .45 - .52 - .38 -
B1: ResRoFA-MT .46 - .48 - .53 - .43 -
B2: 20% subset v2 - - - - .43 - .44 -
B3a: Glove-shared-not-tuned .30 - .34 - .49 - .31 -
B3b: Glove-shared-tuned .29 - .29 - .53 - .33 -
B3c: Glove-shared-tuned with 20%M .35 - .43 - .59 - .43 -
Exp.1.1 .63 .121 .64 .132 .35 .168 .54 .159
Exp.1.2 .61 .081 .65 .120 .40 .142 .58 .131
Exp.2.1 .48 .085 .49 .075 .42 .149 .55 .136
Exp.2.2 .46 .063 .54 .081 .45 .154 .56 .130
Exp.3.1 .69 .097 .70 .115 .48 .125 .65 .137
Exp.3.2 .68 .108 .68 .132 .46 .118 .66 .138
Exp.4.1 .57 .077 .66 .074 .58 .130 .58 .133
Exp.4.2 .58 .091 .66 .092 .59 .132 .59 .145
Table 7: GPT4-Turbo Spearman Rank Correlation Coefficient ($\rho$) with variance.
Model Dataset Fer-Loc Fer-Ins Pado McRae
$\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$
BG: GSD2015 .29 - .42 - .53 - .36 -
B0: NN RF .44 - .45 - .52 - .38 -
B1: ResRoFA-MT .46 - .48 - .53 - .43 -
B2: 20% subset v2 - - - - .43 - .44 -
B3a: Glove-shared-not-tuned .30 - .34 - .49 - .31 -
B3b: Glove-shared-tuned .29 - .29 - .53 - .33 -
B3c: Glove-shared-tuned with 20%M .35 - .43 - .59 - .43 -
Exp.1.1 .51 .046 .56 .105 .40 .145 .64 .158
Exp.1.2 .50 .051 .56 .105 .40 .147 .58 .149
Exp.2.1 .32 .024 .33 .050 .33 .096 .40 .058
Exp.2.2 .26 .027 .37 .061 .33 .114 .37 .066
Exp.3.1 .59 .042 .59 .082 .45 .111 .64 .118
Exp.3.2 .56 .062 .59 .097 .42 .127 .63 .136
Exp.4.1 .56 .030 .53 .056 .42 .104 .44 .061
Exp.4.2 .53 .043 .53 .071 .42 .121 .46 .074
Table 8: GPT4.1 Spearman Rank Correlation Coefficient ($\rho$) with variance.
Appendix C Earlier round of Experiment
A previous round of these experiments has been done on GPT4-Turboand CodeLlama2 (codeLlama2-70B-instruct) (roziere2023code) with a previous version of the prompts we use. The results of the earlier round are shown in Table 9 and Table 10.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .62 .73 .56 .50
Exp.1.2 .62 .74 .54 .48
Exp.2.1 .43 .63 .52 .48
Exp.2.2 .52 .68 .54 .51
Exp.3.1 .77 .67 .28 .33
Exp.3.2 .77 .68 .28 .36
Exp.4.1 .53 .66 .54 .42
Exp.4.2 .57 .68 .53 .45
Table 9: GPT4-Turbo Spearman Rank Correlation coefficient ($\rho$) results (old prompts). All $\rho$’s had p-values $10^{-7}$.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .31 .42 .08 .22
Exp.1.2 .18 .22 -.06 .00
Exp.2.1 .26 .43 .16 .17
Exp.2.2 .21 .19 -.04 -.01
Exp.3.1 .48 .50 .14 .15
Exp.3.2 .21 .37 .16 .10
Exp.4.1 .39 .46 .20 .13
Exp.4.2 .14 .36 -.01 .09
Table 10: CodeLlama2 Spearman Rank Correlation coefficient ($\rho$) results (old prompts). Same models as in Table 9. All $\rho$’s for Fer-Loc and Fer-Ins had p<.003; for Pado and McRae we had p<.84 and p<.94, respectively (very high p-value).
Appendix D Hyperparameters Preliminary Study
Table 11 shows the results of the preliminary study that we conducted previously to choose the most suitable values for temperature and top_p of the models. The study was done on a subset of the Fer-Ins dataset (20 samples $\sim$ 8%). We ran a parameter sweep over temperature values (0.0, 0.5, 0.9) and top_p values (0.5, 0.7, 0.95). Temperature values were chosen near-uniformly across the scale to examine a wide range of the model creativity setting. Higher values correspond to higher randomness (aka "creativity") in the output. High values of top_p limit the randomness in the output to more probable tokens. Since we were more interested in accurate output, and less interested in creative writing for our purposes here, we tested top-p threshold values on the higher half of the scale. As a result of this sweep, we set temperature=0 and top_p=.95.
temp top_p .5 .7 .95
0 .77 .72 .80
.5 .52 .71 .75
.9 .61 .78 .74
Table 11: Preliminary Study for temperature and top_p; each cell represents the respective Spearman rank Correlation Coefficient ($\rho$)
Appendix E Analysis: Why did reasoning have mixed effects?
The achievements of Step-by-Step Prompting in arithmetic, commonsense, and symbolic reasoning motivated us to test its potential in linguistic tasks. The linguistic tasks are similar to these arithmetic tasks in that they both require Step-by-Step Prompting, and having a (virtual or physical) scratchpad for these intermediate steps or calculations, helps get to the final answer. However, different tasks require different types of knowledge and reasoning capability. The thematic fit task requires a semantic reasoning ability and the ability to provide graded, non-binary judgment. Our experiments, conducted with this “scratchpad hypothesis” in mind, did not turn out as expected. As shown in Table 9, the effect of Step-by-Step Prompting shines with the Fer-Loc dataset, but we observe the success of Simple Prompting over Step-by-Step Prompting on other datasets.
Thus, we analyzed 50 random results across all datasets from Exp.3 in our previous attempt to address the question: Can good/bad reasoning (as defined below) positively/negatively affect the model’s performance on thematic fit estimation tasks?
We define the following four terms:
Good Reasoning: The type of reasoning that includes correct facts/properties about the specified role/argument and is logically valid with the given predicate, arguments, and arguments’ role.
Bad Reasoning: The type of reasoning that includes incorrect facts/properties about the specified role/argument and/or is logically invalid with the given predicate, argument, and argument’s role.
Good Scoring: The difference between the human raters’ (normalized) scores and the models’ scores is less than an arbitrary threshold.
Bad Scoring: The difference between the human raters’ scores and the models’ scores is greater than the threshold.
Classifying the reasoning result of each prompt in the sample as good or bad has been done manually. Classifying the thematic fit score result of each prompt in the sample as good or bad has been done according to the following steps:
  1. 1.
    Normalizing the human-based scores by converting them from a 1-7 scale to a 0-1 scale.
  2. 2.
    Comparing the difference between the normalized human-based scores to model-based scores for each test set item in the sample.
  3. 3.
    We experimented with multiple thresholds to classify the scores. This threshold is a heuristic calculated based on the six-interval normalized human scale. The first threshold is less than one interval (10%), the second is 1.5 the interval (25%), and the last one is more than three times the interval (50%). The most reasonable threshold was 25%, which is moderate between a strict threshold (10%) and a loose one (50%).
Figure 2: Effect of Early Bad Reasoning. The example was taken from preliminary experimentation.
Reasoning Type Fer-Loc Fer-Inst Pado McRae
Score Type Num Cat Num Cat Num Cat Num Cat
Prompt 1 Good Reasoning 33:17 39:11 29:21 32:18 15:15 20:10 21:19 23:17
Bad Reasoning 0:0 0:0 0:0 0:0 8:12 8:12 8:2 8:2
Prompt 2 Good Reasoning 33:17 39:11 29:21 32:18 19:20 24:15 28:20 30:18
Bad Reasoning 0:0 0:0 0:0 0:0 4:6 4:6 1:1 1:1
Prompt 3 Good Reasoning 33:17 39:11 29:21 32:18 17:15 22:10 23:18 25:16
Bad Reasoning 0:0 0:0 0:0 0:0 6:12 6:12 6:3 6:3
Overall Good Reasoning 33:17 39:11 29:21 32:18 15:13 20:8 20:18 22:16
Bad Reasoning 0:0 0:0 0:0 0:0 8:14 8:14 9:3 9:3
Table 12: Good-to-bad score ratio across the board. ’Num’ means numeric score, whereas ’Cat’ means categorical score. The overall reasoning is considered ‘good’ if all prompts’ responses were ‘good’. Otherwise, it is considered bad reasoning.
The result of this analysis is shown in Table 12. On Pado and McRae, most cases have a high mass of ‘good scores’ as opposed to ‘bad scores’ with good reasoning. Cases of bad reasoning were fewer and far apart. Also, the good-to-bad score ratio in good reasoning was much higher compared to that ratio in bad reasoning. On the Ferretti sets, there were no bad reasoning cases, but the same good-to-bad score ratio pattern was observed. This pattern is reflected in both numeric and categorical scores, with minor differences.
Returning to our question: Can good/bad reasoning (as defined above) positively/negatively affect the model’s performance on thematic fit estimation tasks?
[…]